# Parsimonious Compactly Supported Covariance Models in the Gauss Hypergeometric Class: Identifiability, Reparameterizations, and Asymptotic Properties

### v2

@@ A.10 Proof of Lemma 1
Let $0<a_{1}<a_{2}$ with $a_{1},a_{2}\in I$. For any data vector $\boldsymbol{Z}_{n}$,
$\frac{\widehat{\sigma}_{n}^{2}(a_{1})}{a_{1}^{1+2\kappa}}-\frac{\widehat{\sigma}_{n}^{2}(a_{2})}{a_{2}^{1+2\kappa}}=\frac{1}{n}\,\boldsymbol{Z}_{n}^{\top}\Big(a_{1}^{-(1+2\kappa)}R_{n}(a_{1})^{-1}-a_{2}^{-(1+2\kappa)}R_{n}(a_{2})^{-1}\Big)\boldsymbol{Z}_{n}.$
This expression is nonnegative for all $\boldsymbol{Z}_{n}$ if and only if
$a_{1}^{-(1+2\kappa)}R_{n}(a_{1})^{-1}-a_{2}^{-(1+2\kappa)}R_{n}(a_{2})^{-1}\succeq 0.$
Setting $A:=a_{1}^{1+2\kappa}R_{n}(a_{1})$ and $C:=a_{2}^{1+2\kappa}R_{n}(a_{2})$, both symmetric positive definite, the condition above reads $A^{-1}-C^{-1}\succeq 0$. Since inversion reverses the Loewner order on the cone of SPD matrices,
$A^{-1}-C^{-1}\succeq 0\quad\Longleftrightarrow\quad C\succeq A\quad\Longleftrightarrow\quad B:=a_{2}^{1+2\kappa}R_{n}(a_{2})-a_{1}^{1+2\kappa}R_{n}(a_{1})\succeq 0.$
The $(i,j)$ entry of $B$ is
$B_{ij}=a_{2}^{1+2\kappa}\,\mathcal{H}_{\kappa,\mu,a_{2},d}(\|s_{i}-s_{j}\|)-a_{1}^{1+2\kappa}\,\mathcal{H}_{\kappa,\mu,a_{1},d}(\|s_{i}-s_{j}\|).$
By Bochner’s theorem, $B\succeq 0$ for every finite collection of locations $\{s_{1},\ldots,s_{n}\}\subset D$ if and only if the associated spectral density difference is nonnegative for all $z\geq 0$:
$f(z):=a_{2}^{1+2\kappa}\,\widehat{\mathcal{H}}_{\kappa,\mu,a_{2},d}(z)-a_{1}^{1+2\kappa}\,\widehat{\mathcal{H}}_{\kappa,\mu,a_{1},d}(z)\geq 0.$
Using the scaling identity $\widehat{\mathcal{H}}_{\kappa,\mu,t,d}(z)=t^{d}\,\widehat{\mathcal{H}}_{\kappa,\mu,1,d}(tz)$ induced by (7), the condition $f(z)\geq 0$ for all $z\geq 0$ is equivalent to monotonicity of
$g(t):=t^{2\kappa+1+d}\,\widehat{\mathcal{H}}_{\kappa,\mu,1,d}(t),\qquad t>0.$
Using (7) with $\delta=\frac{d+1}{2}+\kappa$, $\beta=\delta+\frac{\mu}{2}$, $\gamma=\beta+l$, we have, up to a positive constant,
$g(t)=t^{2\kappa+1+d}\,{}_{1}F_{2}\!\left(\delta;\beta,\gamma;-\frac{t^{2}}{4}\right).$
Substituting the power series representation
$_{1}F_{2}\!\left(\delta;\beta,\gamma;-\frac{t^{2}}{4}\right)=\sum_{k=0}^{\infty}\frac{(\delta)_{k}}{(\beta)_{k}(\gamma)_{k}}\frac{(-t^{2}/4)^{k}}{k!},$
and differentiating term by term (justified by uniform convergence on compact subsets of $(0,\infty)$), one obtains
$g^{\prime}(t)$ $=(2\kappa+1+d)\,t^{2\kappa+d}\sum_{k=0}^{\infty}\frac{(\delta)_{k}}{(\beta)_{k}(\gamma)_{k}}\frac{(-t^{2}/4)^{k}}{k!}+t^{2\kappa+1+d}\sum_{k=1}^{\infty}\frac{(\delta)_{k}}{(\beta)_{k}(\gamma)_{k}}\frac{(-1/4)^{k}\,2k\,t^{2k-1}}{k!}$
$=(2\kappa+1+d)\,t^{2\kappa+d}\sum_{k=0}^{\infty}\frac{(\delta)_{k+1}}{(\delta)_{1}(\beta)_{k}(\gamma)_{k}}\frac{(-t^{2}/4)^{k}}{k!}\cdot\frac{(\delta)_{1}}{\delta}\;+\;\cdots$
Using the identity $(\delta)_{k+1}=\delta\,(\delta+1)_{k}$ and collecting terms, both sums combine into
$g^{\prime}(t)=C\,t^{2\kappa+d}\,{}_{1}F_{2}\!\left(\delta+1;\beta,\gamma;-\frac{t^{2}}{4}\right),\qquad t>0,$
where $C=2(\kappa+1+d/2)\cdot\delta/(\beta\gamma)\cdot\beta\gamma>0$. Hence $g$ is nondecreasing on $(0,\infty)$ if and only if
$_{1}F_{2}\!\left(\delta+1;\beta,\gamma;-\frac{t^{2}}{4}\right)\geq 0\qquad\forall\,t>0.$
Fix $l=\frac{d}{2}+\kappa$. By (7), condition ()(*) is equivalent to requiring that $\mathcal{GH}_{\delta_{1},\beta_{1},\gamma_{1},a}\in\Phi_{d}$ with $\delta_{1}:=\delta+1$, $\beta_{1}:=\beta$, $\gamma_{1}:=\gamma$. Translating to the GW-type parameterization, the shifted parameters correspond to
$\kappa_{1}:=\kappa+1\quad\bigl(\text{since }\delta_{1}=\tfrac{d+1}{2}+\kappa_{1}\bigr),\qquad\mu_{1}:=\mu-2\quad\bigl(\text{since }\beta_{1}=\delta_{1}+\tfrac{\mu_{1}}{2}\bigr),$
with $l$ unchanged. Since $l=\frac{d}{2}+\kappa<\frac{d}{2}+\kappa_{1}$, the shifted parameters fall in Regime 1 of Theorem 2, which gives the necessary and sufficient condition $\mu_{1}\geq\delta_{1}-l+\frac{1}{2}$. Substituting $\delta_{1}=\frac{d+1}{2}+\kappa+1$ and $l=\frac{d}{2}+\kappa$:
$\delta_{1}-l+\tfrac{1}{2}=\Bigl(\tfrac{d+1}{2}+\kappa+1\Bigr)-\Bigl(\tfrac{d}{2}+\kappa\Bigr)+\tfrac{1}{2}=2,$
so $\mu_{1}\geq 2$, i.e. $\mu\geq 4$. This completes the proof. ∎

### v3

@@ A.10 Proof of Theorem 11
Fix $\mu>1+\frac{d}{2}$ and $I=[a_{L},a_{U}]$. For $a\in I$ set
$G_{n}(a):=\frac{a^{2\kappa+1}}{F(\kappa,\mu,d)}\,R_{n}(\mu,a),\qquad\hat{\theta}_{n}(\mu,a)=\frac{1}{n}\,\mathbf{Z}_{n}^{\top}G_{n}(a)^{-1}\mathbf{Z}_{n}.$
Let $G_{0n}:=G_{n}(a_{0})$ and define the vector
$\mathbf{Y}_{n}:=\theta_{0}^{-1/2}\,G_{0n}^{-1/2}\mathbf{Z}_{n}\sim\mathcal{N}(0,I_{n}),\qquad\theta_{0}=\sigma_{0}^{2}F(\kappa,\mu,d)/a_{0}^{2\kappa+1}.$
Then
$\frac{\hat{\theta}_{n}(\mu,a)}{\theta_{0}}=\frac{1}{n}\,\mathbf{Y}_{n}^{\top}M_{n}(a)\mathbf{Y}_{n},\qquad M_{n}(a):=G_{0n}^{1/2}G_{n}(a)^{-1}G_{0n}^{1/2}.$ (43)
Fix $a^{\star}\in I$ and set $\Delta_{n}(a):=M_{n}(a)-M_{n}(a^{\star})$. The decomposition
$\frac{\sqrt{n}\,\bigl(\hat{\theta}_{n}(\mu,a)-\hat{\theta}_{n}(\mu,a^{\star})\bigr)}{\theta_{0}}=\underbrace{\frac{1}{\sqrt{n}}\operatorname{tr}(\Delta_{n}(a))}_{\mathrm{(I)}}+\underbrace{\frac{1}{\sqrt{n}}\bigl(\mathbf{Y}_{n}^{\top}\Delta_{n}(a)\mathbf{Y}_{n}-\operatorname{tr}(\Delta_{n}(a))\bigr)}_{\mathrm{(II)}}$ (44)
will be used to prove (ii), and then (i) follows from (ii) by taking $a^{\star}=a_{0}$.
Hereafter, we use $\|A\|_{\mathrm{HS}}^{2}:=\mathrm{tr}(A^{*}A)$ for Hilbert–Schmidt operators $A$, where $A^{*}$ denotes the adjoint operator. For matrices $B$, we use $\|B\|_{F}^{2}:=\mathrm{tr}(B^{\top}B)$.
Let $g_{a}$ denote the isotropic spectral density of the stationary covariance function $G(\cdot\,;a):=\frac{a^{2\kappa+1}}{F(\kappa,\mu,d)}\mathcal{H}_{\kappa,\mu,a,d}(\cdot)$. By Lemma 1 specialised to the $\mathcal{H}$ model, the high-frequency expansion of $g_{a}$ has the form
$g_{a}(z)=C_{\kappa,\mu,d}\,z^{-2\delta}\bigl(1+\mathcal{O}(z^{-2})\bigr)+\mathcal{O}(z^{1-\mu-2\delta}),\qquad z\to\infty,$
with $\delta=\frac{d+1}{2}+\kappa$ and with $\mathcal{O}$-constants uniform over $a\in I$ (the leading term is independent of $a$ because of the prefactor $a^{2\kappa+1}$). Hence, for any fixed $c>0$,
$\sup_{a\in I}\int_{c}^{\infty}z^{d-1}\Bigl(\frac{g_{a}(z)-g_{a_{0}}(z)}{g_{a_{0}}(z)}\Bigr)^{2}\,dz<\infty,$ (45)
since $\mu>1+d/2$ and $d\in\{1,2,3\}$ make the tail square-integrable. By the fixed-domain equivalence criterion of Stein (2004) and its connection to covariance-ratio operators (see Skorokhod and Yadrenko (1973, Assertion B)), (45) implies that the operator
$\mathcal{M}(a):=\mathcal{G}(a_{0})^{1/2}\mathcal{G}(a)^{-1}\mathcal{G}(a_{0})^{1/2}-\mathcal{I}\quad\text{on }L^{2}(D)$
is Hilbert–Schmidt, with $\sup_{a\in I}\|\mathcal{M}(a)\|_{\mathrm{HS}}<\infty$. Let $\Pi_{n}$ denote the orthogonal projector onto the finite-dimensional subspace generated by the observations $\{Z(\mathbf{s}_{1}),\ldots,Z(\mathbf{s}_{n})\}$ (identified with its image in $L^{2}(D)$ through the covariance-operator embedding). Then $M_{n}(a)-I_{n}$ represents the restriction of $\Pi_{n}\mathcal{M}(a)\Pi_{n}$ to that subspace, hence
$\sup_{a\in I}\|M_{n}(a)-I_{n}\|_{F}^{2}=\sup_{a\in I}\|\Pi_{n}\mathcal{M}(a)\Pi_{n}\|_{\mathrm{HS}}^{2}\leq\sup_{a\in I}\|\mathcal{M}(a)\|_{\mathrm{HS}}^{2}=:C_{\mathrm{HS}}<\infty,\qquad\forall n.$ (46)
In particular,
$\sup_{a\in I}\|M_{n}(a)\|$ $\leq 1+\sup_{a\in I}\|M_{n}(a)-I_{n}\|\leq 1+\sup_{a\in I}\|M_{n}(a)-I_{n}\|_{F}$ (47)
$\leq 1+\sqrt{C_{\mathrm{HS}}},\qquad\sup_{a\in I}\|\Delta_{n}(a)\|_{F}^{2}\leq 4C_{\mathrm{HS}},\qquad\forall n.$
Let $p:=\mu+\frac{d-1}{2}+2\kappa$ and $\alpha:=\min(1,p)$. Since $\mu>1+\frac{d}{2}$ and $\kappa>-1/2$, one has $\alpha>1/2$. From the explicit form of $\mathcal{H}_{\kappa,\mu,a,d}$ and its compact support in $[0,a]$, the map $a\mapsto\mathcal{H}_{\kappa,\mu,a,d}(r)$ is uniformly $\alpha$-Hölder on $I$, namely
$\sup_{r\geq 0}\bigl|\mathcal{H}_{\kappa,\mu,a,d}(r)-\mathcal{H}_{\kappa,\mu,a^{\prime},d}(r)\bigr|\leq C_{\alpha}|a-a^{\prime}|^{\alpha},\qquad a,a^{\prime}\in I,$ (48)
where the exponent $\alpha=\min(1,p)$ reflects the boundary behaviour at $r=a$. Moreover, using (48), continuity in $(a,z)$ on compact sets, positivity of $g_{a_{0}}$, and the common high-frequency expansion above, one obtains that for any fixed $c>0$,
$\int_{c}^{\infty}z^{d-1}\Bigl(\frac{g_{a}(z)-g_{a^{\prime}}(z)}{g_{a_{0}}(z)}\Bigr)^{2}dz\leq C|a-a^{\prime}|^{2\alpha},\qquad a,a^{\prime}\in I,$
with $C$ independent of $(a,a^{\prime})$. Consequently, by the same covariance-ratio/operator correspondence (Stein, 2004, Appendix A) (see also Skorokhod and Yadrenko (1973, Assertion B)),
$\|\mathcal{M}(a)-\mathcal{M}(a^{\prime})\|_{\mathrm{HS}}\leq C|a-a^{\prime}|^{\alpha},\qquad a,a^{\prime}\in I.$
Passing to the discretised matrices and using again that orthogonal projections do not increase HS norms yields the Frobenius increment bound
$\|\Delta_{n}(a)-\Delta_{n}(a^{\prime})\|_{F}\leq C|a-a^{\prime}|^{\alpha},\qquad a,a^{\prime}\in I,\ \forall n.$ (49)
By Hanson–Wright (Hanson and Wright, 1971; Rudelson and Vershynin, 2013), for any symmetric $A$ and $\mathbf{Y}\sim\mathcal{N}(0,I_{n})$,
$P\!\Bigl(\bigl|\mathbf{Y}^{\top}A\mathbf{Y}-\operatorname{tr}(A)\bigr|>t\Bigr)\leq 2\exp\!\Bigl(-c\min\!\Bigl\{\frac{t^{2}}{\|A\|_{F}^{2}},\frac{t}{\|A\|}\Bigr\}\Bigr).$
With $A=\Delta_{n}(a)$ and $t=\varepsilon\sqrt{n}$, using (47) and $\|A\|\leq\|A\|_{F}$,
$P\!\Bigl(\Bigl|\frac{1}{\sqrt{n}}\bigl(\mathbf{Y}_{n}^{\top}\Delta_{n}(a)\mathbf{Y}_{n}-\mathrm{tr}\Delta_{n}(a)\bigr)\Bigr|>\varepsilon\Bigr)\leq 2e^{-c^{\prime}n},$
uniformly in $a\in I$. Let $\delta_{n}:=n^{-4}$ and let $\mathcal{G}_{n}\subset I$ be a regular grid of mesh $\delta_{n}$ (so $|\mathcal{G}_{n}|=\mathcal{O}(n^{4})$). A union bound gives
$\sup_{a\in\mathcal{G}_{n}}\Bigl|\frac{1}{\sqrt{n}}\bigl(\mathbf{Y}_{n}^{\top}\Delta_{n}(a)\mathbf{Y}_{n}-\mathrm{tr}\Delta_{n}(a)\bigr)\Bigr|=o_{p}(1).$ (50)
For arbitrary $a\in I$, choose $a^{(n)}\in\mathcal{G}_{n}$ with $|a-a^{(n)}|\leq\delta_{n}$. Then by (49) and $\|\mathbf{Y}_{n}\|^{2}=O_{p}(n)$,
$\frac{1}{\sqrt{n}}\bigl|\mathbf{Y}_{n}^{\top}(\Delta_{n}(a)-\Delta_{n}(a^{(n)}))\mathbf{Y}_{n}\bigr|$ $\leq\frac{\|\Delta_{n}(a)-\Delta_{n}(a^{(n)})\|_{F}}{\sqrt{n}}\,\|\mathbf{Y}_{n}\|^{2}$
$\leq C\delta_{n}^{\alpha}\,O_{p}(\sqrt{n})=O_{p}(n^{1/2-4\alpha})=o_{p}(1).$
since $\alpha>1/2$. Combining with (50) yields
$\sup_{a\in I}\Bigl|\frac{1}{\sqrt{n}}\bigl(\mathbf{Y}_{n}^{\top}\Delta_{n}(a)\mathbf{Y}_{n}-\mathrm{tr}\Delta_{n}(a)\bigr)\Bigr|=o_{p}(1).$ (51)
Now, define
$r_{n}(a):=\frac{1}{n}\mathrm{tr}(M_{n}(a))-1=\frac{E[\hat{\theta}_{n}(\mu,a)]}{\theta_{0}}-1.$
Then $\frac{1}{\sqrt{n}}\mathrm{tr}(\Delta_{n}(a))=\sqrt{n}\bigl(r_{n}(a)-r_{n}(a^{\star})\bigr)$. First, by (49) and Cauchy–Schwarz,
$\sqrt{n}\,|r_{n}(a)-r_{n}(a^{\prime})|=\frac{|\mathrm{tr}(M_{n}(a)-M_{n}(a^{\prime}))|}{\sqrt{n}}\leq\|M_{n}(a)-M_{n}(a^{\prime})\|_{F}\leq C|a-a^{\prime}|^{\alpha},$
uniformly in $n$, so $\{\sqrt{n}\,r_{n}(\cdot)\}_{n\geq 1}$ is equicontinuous on $I$. Second, for each fixed $a$, Theorem 10(ii) gives $\sqrt{n}(\hat{\theta}_{n}(\mu,a)-\theta_{0})\Rightarrow\mathcal{N}(0,2\theta_{0}^{2})$. Moreover, for Gaussian quadratic forms,
$\rm Var\!\Big(\mathbf{Y}_{n}^{\top}M_{n}(a)\mathbf{Y}_{n}-\operatorname{tr}(M_{n}(a))\Big)=2\,\operatorname{tr}(M_{n}(a)^{2})\leq 2\,\|M_{n}(a)\|_{F}^{2},$
so (46) yields a uniform second-moment bound and hence uniform integrability. Therefore
[…]
By equicontinuity and compactness of $I$ (finite-cover argument),
$\sup_{a\in I}\frac{1}{\sqrt{n}}|\mathrm{tr}(\Delta_{n}(a))|=\sup_{a\in I}\sqrt{n}\,|r_{n}(a)-r_{n}(a^{\star})|\leq 2\sup_{a\in I}\sqrt{n}\,|r_{n}(a)|\longrightarrow 0.$ (52)
Finally, combining (44), (51), and (52) gives
$\sup_{a\in I}\sqrt{n}\,\bigl|\hat{\theta}_{n}(\mu,a)-\hat{\theta}_{n}(\mu,a^{\star})\bigr|=o_{p}(1),$
which proves (ii).
To show (i), take $a^{\star}=a_{0}$ in (ii) to get
$\sup_{a\in I}|\hat{\theta}_{n}(\mu,a)-\theta_{0}|\leq\sup_{a\in I}|\hat{\theta}_{n}(\mu,a)-\hat{\theta}_{n}(\mu,a_{0})|+|\hat{\theta}_{n}(\mu,a_{0})-\theta_{0}|.$
The first term is $o_{p}(1)$ by (ii), and the second converges to 0 almost surely by Theorem 10(i). Hence $\sup_{a\in I}|\hat{\theta}_{n}(\mu,a)-\theta_{0}|\to 0$ in probability, proving (i). ∎
A.11 Proof of Theorem 12
The proof of Theorem 12 requires a uniform stochastic control of the plug-in microergodic estimator over the compact parameter set $J\times I$, since the ML estimators $(\hat{\mu}_{n},\hat{a}_{n})$ are random. The next lemma extends Lemma 2 to the case where $\mu$ is unknown and varies in $J$.
Lemma 3.
Assume the setting of Theorem 10, with $(\sigma_{0}^{2},\mu_{0},a_{0})\in(0,\infty)\times J\times I$ and $\mu_{0}>1+\frac{d}{2}$, where $J=[\mu_{L},\mu_{U}]$ and $I=[a_{L},a_{U}]$ with $\mu_{L}>1+\frac{d}{2}$. Let
$\hat{\theta}_{n}(\mu,a)=\frac{\hat{\sigma}_{n}^{2}(\mu,a)\,F(\kappa,\mu,d)}{a^{2\kappa+1}},\qquad\hat{\sigma}_{n}^{2}(\mu,a)=\frac{\mathbf{Z}_{n}^{\top}R_{n}(\mu,a)^{-1}\mathbf{Z}_{n}}{n},$
and $\theta_{0}=\sigma_{0}^{2}F(\kappa,\mu_{0},d)/a_{0}^{2\kappa+1}$. Then, as $n\to\infty$,
  1. (i)
    $\sup_{(\mu,a)\in J\times I}\big|\hat{\theta}_{n}(\mu,a)-\theta_{0}\big|\xrightarrow{p}0;$
  2. (ii)
    for any fixed $(\mu^{\star},a^{\star})\in J\times I$,
    $\sup_{(\mu,a)\in J\times I}\sqrt{n}\,\big|\hat{\theta}_{n}(\mu,a)-\hat{\theta}_{n}(\mu^{\star},a^{\star})\big|=o_{p}(1).$
Proof.
The argument follows closely the proof of Lemma 2. We keep the same notation and decomposition, replacing $a\in I$ by $(\mu,a)\in J\times I$, and we record only the modifications needed to obtain uniformity over the two-parameter compact set.
All bounds that in Lemma 2 were uniform over $a\in I$ remain uniform over $(\mu,a)\in J\times I$, because $J\times I$ is compact and $\mu_{L}>1+\frac{d}{2}$. In particular, the high-frequency expansion of the normalised spectral density and Stein’s integrability condition hold uniformly on $J\times I$, yielding uniform Hilbert–Schmidt bounds for the covariance-ratio operators and hence
$\sup_{(\mu,a)\in J\times I}\|M_{n}(\mu,a)-I_{n}\|_{F}^{2}\leq C,\qquad\sup_{(\mu,a)\in J\times I}\|M_{n}(\mu,a)\|\leq 1+\sqrt{C},$
for some finite constant $C$ and all $n$, where $M_{n}(\mu,a)$ is defined as in Lemma 2 with $G_{n}(\mu,a):=\frac{a^{2\kappa+1}}{F(\kappa,\mu,d)}\,R_{n}(\mu,a)$. Moreover, on $J\times I$ the map $\mu\mapsto\mathcal{H}_{\kappa,\mu,a,d}(r)$ is $C^{1}$ and thus uniformly Lipschitz in $\mu$, while the map $a\mapsto\mathcal{H}_{\kappa,\mu,a,d}(r)$ is uniformly $\alpha$–Hölder in $a$ with $\alpha>1/2$, implying the two-parameter increment bound
$\|\Delta_{n}(\mu,a)-\Delta_{n}(\mu^{\prime},a^{\prime})\|_{F}\leq C\bigl(|\mu-\mu^{\prime}|+|a-a^{\prime}|^{\alpha}\bigr),\qquad(\mu,a),(\mu^{\prime},a^{\prime})\in J\times I,$
where $\Delta_{n}(\mu,a)$ denotes the corresponding increment matrix.
Finally, the Hanson–Wright and grid argument is applied on a two-dimensional grid of mesh $\delta_{n}=n^{-4}$ in $J\times I$ (of cardinality $\mathcal{O}(n^{8})$), and the remainder of the proof proceeds exactly as in Lemma 2, yielding (ii). Statement (i) then follows by taking $(\mu^{\star},a^{\star})=(\mu_{0},a_{0})$ and using Theorem 10(i). ∎
Proof of Theorem 12 .
By definition, $(\hat{\mu}_{n},\hat{a}_{n})\in J\times I$ for all $n$.
Proof of (i). Since $(\hat{\mu}_{n},\hat{a}_{n})\in J\times I$,
$\big|\hat{\theta}_{n}(\hat{\mu}_{n},\hat{a}_{n})-\theta_{0}\big|\leq\sup_{(\mu,a)\in J\times I}\big|\hat{\theta}_{n}(\mu,a)-\theta_{0}\big|.$
By Lemma 3(i), the right-hand side converges to 0 in probability, hence $\hat{\theta}_{n}(\hat{\mu}_{n},\hat{a}_{n})\xrightarrow{p}\theta_{0}$.
Proof of (ii). Fix the nonrandom point $(\mu^{\star},a^{\star})=(\mu_{0},a_{0})\in J\times I$. Then
$\sqrt{n}\big(\hat{\theta}_{n}(\hat{\mu}_{n},\hat{a}_{n})-\theta_{0}\big)$ $=\sqrt{n}\big(\hat{\theta}_{n}(\mu_{0},a_{0})-\theta_{0}\big)$
$\quad+\sqrt{n}\big(\hat{\theta}_{n}(\hat{\mu}_{n},\hat{a}_{n})-\hat{\theta}_{n}(\mu_{0},a_{0})\big).$
By Lemma 3(ii), the second term is $o_{p}(1)$. By Theorem 10(ii) (applied with $\mu=\mu_{0}$ and $a=a_{0}$ fixed),
$\sqrt{n}\big(\hat{\theta}_{n}(\mu_{0},a_{0})-\theta_{0}\big)\xrightarrow{\mathcal{D}}\mathcal{N}(0,2\theta_{0}^{2}).$
Therefore, Slutsky’s theorem yields
$\sqrt{n}\big(\hat{\theta}_{n}(\hat{\mu}_{n},\hat{a}_{n})-\theta_{0}\big)\xrightarrow{\mathcal{D}}\mathcal{N}(0,2\theta_{0}^{2}),$
which proves (ii). ∎